# A new web of dualities from Majorana Fermions

### v2

@@ I Introduction
Dualities are a central theme in modern theoretical physics, providing complementary descriptions of the same physical phenomena and often uncovering deep structural relations between seemingly distinct theories. In quantum field theory, two models are said to be dual when their physical observables coincide, even though their microscopic degrees of freedom and Lagrangian formulations may differ drastically [Polchinski:2014mva]. Among the earliest and most influential examples is two-dimensional bosonization, first established by Coleman [Coleman:1974bu], which identifies a massless Dirac fermion with a compactified scalar field through a precise mapping of operators and couplings.
@@
In three dimensions, bosonization manifests as an infrared (IR) duality, emerging at the long-distance fixed points of distinct microscopic theories. Two QFTs are IR-dual if their renormalization-group flows terminate at the same conformal field theory (CFT). Such dualities have attracted renewed interest [Karch:2016sxi, Seiberg:2016gmd, Witczak-Krempa:2023lgm, Metlitski:2016dht, Aharony:2016jvv], owing both to their connection with strongly coupled CFTs and to their relevance in condensed-matter systems near quantum criticality [Seiberg:2016gmd].  [^1] See also [Amoretti:2013xya], for the emergence of 3+1D Weyl fermions from bosonic salar tensor dualities.
@@
A systematic approach to generating new IR dualities was proposed in [Witczak-Krempa:2023lgm]. Starting from the conjectured three-dimensional Dirac bosonization [Karch:2016sxi, Seiberg:2016gmd], relevant operators are identified on both sides of the duality and used to deform the theories away from the critical point. When these deformations are promoted to dynamical scalar fields, the resulting renormalization-group flows lead to new interacting fixed points conjectured to be dual. The validity of these conjectures can be tested through the matching of global symmetries, low-energy degrees of freedom, and anomalies.
@@
While the construction in [Witczak-Krempa:2023lgm] involves Dirac fermions, a parallel bosonization for Majorana fermions has been proposed in [Metlitski:2016dht, Aharony:2016jvv]. It relates a single neutral Majorana spinor to an $SO(N)_{1}$ Chern–Simons theory coupled to an $N$-component real scalar field. At criticality, both theories describe the transition between a trivial insulator and a $p_{x}+ip_{y}$ topological superconductor, whose edge hosts a chiral Majorana mode [Stone:2003wbl, Gnezdilov:2015twa].
@@
The purpose of this work is to incorporate the Majorana bosonization into the existing duality web and to demonstrate that it naturally generates new classes of non-Abelian boson–boson equivalences. Starting from the two-flavour case, we derive a novel correspondence between an $SO(N)_{1}$ Chern–Simons theory coupled to two real scalars and a $U(1)_{1}$ theory coupled to a single complex scalar, with consistent boundary conditions ensuring chiral-central-charge matching. We then generalize the construction to an arbitrary even number of Majorana flavours, obtaining a hierarchy of dualities relating $SO(N)_{1}$ gauge theories to multiple copies of Abelian Chern–Simons–Higgs models. Furthermore, we study pairing-type deformations of the seed duality and show that they induce controlled Higgs flows $SO(N)\!\to\!SO(N{-}1)$, reproducing the expected thermal Hall responses in the IR. Finally, we propose a broader family of $SO(N~)SO(\tilde{N})$ equivalences that emerge when several Majorana bosonizations are combined, extending the known web of three-dimensional dualities beyond the Dirac sector.
II Generating new dualities from the Dirac seed
@@
An essential ingredient in our construction is the method developed in [Witczak-Krempa:2023lgm]. To illustrate its main features, we begin with the $(2+1)$-dimensional seed duality conjectured in [Karch:2016sxi, Seiberg:2016gmd]:
@@
As noted in [Witczak-Krempa:2023lgm], deforming the two sides by mass terms allows one to probe and match their distinct phases. On the bosonic side, a mass deformation
@@
must have the opposite sign [Seiberg:2016gmd, Polyakov:1988md, Ferreiros:2018ohl],
@@
As usual for mass deformations, these operators are relevant and drive the two theories away from the critical point (1). Identifying pairs of relevant operators that correspond across the duality is the central idea of the method in [Witczak-Krempa:2023lgm]: by deforming both sides with the same operator, one moves along a shared RG trajectory and reaches new IR fixed points that can be conjectured to be dual as well.
@@
III Bosonization of a Majorana fermion
Our main goal is to incorporate Majorana fermions into this framework and to explore the resulting web of dualities. In the previous section we began with a massless theory and discussed how the mass deformation links the deformed theories; for the Majorana case let us begin with the full (conjectured) massive correspondence. The bosonization of a single neutral fermion was first proposed in [Metlitski:2016dht, Aharony:2016jvv], which relates the free theory of a Majorana spinor $\xi$,
@@ III Bosonization of a Majorana fermion
$i\bar{\xi}\cancel{\partial}\xi\;\longleftrightarrow\;|(\partial-ia)\vec{\phi}|^{2}-\lambda(\vec{\phi}^{2})^{2}$
[…]
with the associated identification of relevant operators [Jian:2018amu],
@@
$S_{\phi}^{N_{f}}[\vec{\phi}_{i},a]=\\ \int_{M}d^{3}x\left[\sum_{i=1}^{N_{f}}|D_{a}\vec{\phi}_{i}|^{2}-r\,\text{Tr}M-u(\text{Tr}M)^{2}-v\,\text{Tr}M^{2}\right]\\ +\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,$, (12)
@@
IV Dirac Fermion in terms of Majorana fermions
@@
with each $\xi_{i}$ satisfying the Majorana condition [Tong:QFT]
@@
$\int_{\mathcal{M}}d^{3}x\,\bar{\psi}(i\cancel{\partial}-m)\psi=\\ \int_{\mathcal{M}}d^{3}x\,\big[\bar{\xi}_{1}(i\cancel{\partial}-m)\xi_{1}+\bar{\xi}_{2}(i\cancel{\partial}-m)\xi_{2}\big]\\ +\int_{\partial\mathcal{M}}d^{2}x\,\bar{\xi}_{1}\gamma^{\mu}\xi_{2}\,n_{\mu}\,$, (17)
@@
The boundary $\partial\mathcal{M}$ is $(1+1)$-dimensional, where chirality can be defined as in $(3+1)$ dimensions. Defining $\gamma^{5}=\gamma^{0}\gamma^{1}$}, which anticommutes with all $\gamma^{\mu}$, we can consider chiral boundary modes (eigenstates of $\gamma^{5}$) for the two Majorana flavours. These Majorana–Weyl fermions exist only in $(8k+2)$ spacetime dimensions [Polchinski:1998rr]. On the boundary we define
@@
This result is consistent with the Abelian Chern–Simons description of a Dirac fermion at low energies, which carries $c_{-}=1$} [Tong:2016kpv, Nawata:2022ywk]. As we will emphasize later, the non-trivial boundary structure will play a key role in matching the edge modes across dualities.
Finally, we note that these boundary conditions emerge naturally from the variational principle used to derive the equations of motion. Although boundary conditions are often imposed externally, this approach allows the theory itself to determine consistent constraints [Amoretti:2014iza, Amoretti:2014kba, Amoretti:2013nv].
V Two Majorana flavours
@@
$\mathcal{L}_{\phi}[\vec{\phi}_{1},\vec{\phi}_{2},a]=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}\\ -\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,$, (23)
@@
To analyze the phase structure, we follow the auxiliary-scalar method of [Witczak-Krempa:2023lgm]. Using the operator correspondence from [Metlitski:2016dht, Aharony:2016jvv, Jian:2018amu],
@@
$\mathcal{L}_{\xi}[\xi_{1},\xi_{2}]-\varphi\sum_{i=1}^{2}\bar{\xi}_{i}\xi_{i}-v(\varphi)$
$\;\longleftrightarrow\;\mathcal{L}_{\phi}[\vec{\phi}_{1},\vec{\phi}_{2},a]+\varphi\sum_{i=1}^{2}\vec{\phi}_{i}^{2}-v(\varphi)\,$, (26)
@@
The sign of $r$ determines the two possible regimes. For r>0{r>0}, $\varphi$ is heavy and can be integrated out, restoring the correspondence (24). For r<0{r<0}, $\varphi$ acquires a nonzero expectation value and induces mass terms for both fermions and scalars. On the bosonic side, the resulting quadratic term combines with $\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})$ to reproduce the full potential $V(\vec{\phi}_{1},\vec{\phi}_{2})$ in (13). Depending on the sign of $\langle\varphi\rangle\equiv\varphi_{0}$}, the two theories describe either a trivial insulator or a topological phase with chiral central charge $c_{-}=1$}. These phases coincide with those found in [Witczak-Krempa:2023lgm], as expected from the Dirac decomposition (17).
@@
$\sum_{i=1}^{2}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}$
$\;\longleftrightarrow\;|D_{b}z|^{2}-|z|^{4}+\text{CS}[b]_{1}\,$, (27)
@@
Again, the edge theories differ by one unit of chiral central charge ($c_{-}=1$} on the left, $c_{-}=0$} on the right), reconciled by the same boundary conditions preserving chirality. This transition connects two gapped phases with thermal Hall conductivities 0 and $\pm\tfrac{\pi}{6}T$ at low temperatures [Kapustin:2019zcq, Zhang:2023hgq].
It is illuminating to reinterpret these boundary conditions in the framework of (28). Following [Metlitski:2016dht, Aharony:2016jvv], the Majorana fermion can be identified with the monopole operator of the gauge field, consistent with viewing this bosonization as a relativistic form of flux attachment. For two flavours, this follows from the equation of motion for the temporal component of the gauge field $a^{i,0}$ derived from (12):
@@
where $\mathscr{M}_{a}$ is the monopole operator of the gauge field [Turner:2019wnh]. The boundary condition (21) then translates to
@@
On the right-hand side of (28), the analogous identification involves the Dirac fermion [Witczak-Krempa:2023lgm]:
@@
VI Even number of Majorana fermions
@@
$\mathcal{L}_{\phi}[\vec{\phi}_{i},a]$ $=$ $\sum_{i=1}^{2N_{f}}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\ldots,\vec{\phi}_{2N_{f}})$ (32)
$+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,$,
@@
VII Ising–Gross–Neveu
Following Ref. [Witczak-Krempa:2023lgm], a theory of two Dirac fermions with staggered mass terms exhibits a transition between a trivial insulator and a quantum spin Hall insulator [Cho:2010rk]. Once again, the use of Majorana variables provides a natural framework for constructing infrared dualities among purely bosonic theories.
@@ VII Ising–Gross–Neveu
$\mathcal{L}_{\phi}[\vec{\phi}_{i},a,\tilde{a}]=\sum_{i=1}^{2}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}+\sum_{i=3}^{4}|D_{\tilde{a}}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{3},\vec{\phi}_{4})+\text{CS}_{SO(\tilde{N})}[\tilde{a}]_{1}+(N+\tilde{N})\text{CS}_{g}\,.$ (35)
@@
and analogously for $SO(2)_{\tilde{\phi}}$ acting on $\tilde{\Phi}=(\vec{\phi}_{3},\vec{\phi}_{4})^{\top}$. These are genuine global symmetries, distinct from the $U(1)^{2}$ gauge symmetry acting on the $z_{i}$. On the $U(1)$ side, the corresponding topological symmetries are $U(1)_{T}^{2}$, under which monopole operators carry charge [Turner:2019wnh].
VIII Pairing transition
Another transition discussed in Ref. [Witczak-Krempa:2023lgm] involves a superconducting deformation of the Dirac seed duality [Karch:2016sxi]:

### v3

@@ 1 Introduction
Dualities are a central theme in modern theoretical physics, providing complementary descriptions of the same physical phenomena and often uncovering deep structural relations between seemingly distinct theories. In quantum field theory, two models are said to be dual when their physical observables coincide, even though their microscopic degrees of freedom and Lagrangian formulations may differ drastically [17]. Among the earliest and most influential examples is two-dimensional bosonization, first established by Coleman [7], which identifies a massless Dirac fermion with a compactified scalar field through a precise mapping of operators and couplings.
@@
In three dimensions, bosonization manifests as an infrared (IR) duality, emerging at the long-distance fixed points of distinct microscopic theories. Two QFTs are IR-dual if their renormalization-group flows terminate at the same conformal field theory (CFT). Such dualities have attracted renewed interest [12, 19, 24, 14, 1], owing both to their connection with strongly coupled CFTs and to their relevance in condensed-matter systems near quantum criticality [19] [^1] See also [3], for the emergence of 3+1D Weyl fermions from bosonic salar tensor dualities..
@@
A systematic approach to generating new IR dualities was proposed in [24]. Starting from the conjectured three-dimensional Dirac bosonization [12, 19], relevant operators are identified on both sides of the duality and used to deform the theories away from the critical point. When these deformations are promoted to dynamical scalar fields, the resulting renormalization-group flows lead to new interacting fixed points conjectured to be dual. The validity of these conjectures can be tested through the matching of global symmetries, low-energy degrees of freedom, and anomalies.
@@
While the construction in [24] involves Dirac fermions, a parallel bosonization for Majorana fermions has been proposed in [14, 1]. It relates a single neutral Majorana spinor to an $SO(N)_{1}$ Chern–Simons theory coupled to an $N$-component real scalar field. At criticality, both theories describe the transition between a trivial insulator and a $p_{x}+ip_{y}$ topological superconductor, whose edge hosts a chiral Majorana mode [20, 9].
@@
The purpose of this work is to incorporate the Majorana bosonization into the existing duality web and to demonstrate that it naturally generates new classes of non-Abelian boson–boson equivalences. Starting from the two-flavour case, we derive a novel correspondence between an $SO(N)_{1}$ Chern–Simons theory coupled to two real scalars and a $U(1)_{1}$ theory coupled to a single complex scalar, with consistent boundary conditions ensuring chiral-central-charge matching. We then generalize the construction to an arbitrary even number of Majorana flavours, obtaining a hierarchy of dualities relating $SO(N)_{1}$ gauge theories to multiple copies of Abelian Chern–Simons–Higgs models. Finally, we study pairing-type deformations of the seed duality and show that they induce controlled Higgs flows $SO(N)\!\to\!SO(N{-}1)$, reproducing the expected thermal Hall responses in the IR.
2 Generating new dualities from the Dirac seed
@@
An essential ingredient in our construction is the method developed in [24]. To illustrate its main features, we begin with the $(2+1)$-dimensional seed duality conjectured in [12, 19]:
@@
As noted in [24], deforming the two sides by mass terms allows one to probe and match their distinct phases. On the bosonic side, a mass deformation
@@
must have the opposite sign [19, 18, 8],
@@
As usual for mass deformations, these operators are relevant and drive the two theories away from the critical point (1). Identifying pairs of relevant operators that correspond across the duality is the central idea of the method in [24]: by deforming both sides with the same operator, one moves along a shared RG trajectory and reaches new IR fixed points that can be conjectured to be dual as well.
@@
3 Bosonization of a Majorana fermion
Our main goal is to incorporate Majorana fermions into this framework and to explore the resulting web of dualities. In the previous section we began with a massless theory and discussed how the mass deformation links the deformed theories; for the Majorana case let us begin with the full (conjectured) massive correspondence. The bosonization of a single neutral fermion was first proposed in [14, 1], which relates the free theory of a Majorana spinor $\xi$,
@@ 3 Bosonization of a Majorana fermion
$i\bar{\xi}\cancel{\partial}\xi\;\longleftrightarrow\;|(\partial-ia)\vec{\phi}|^{2}-\lambda(\vec{\phi}^{2})^{2}+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g},$ (9)
with the associated identification of relevant operators [10],
@@
$S_{\phi}^{N_{f}}[\vec{\phi}_{i},a]=\int_{M}d^{3}x\left[\sum_{i=1}^{N_{f}}|D_{a}\vec{\phi}_{i}|^{2}-r\,\text{Tr}M-u(\text{Tr}M)^{2}-v\,\text{Tr}M^{2}\right]\\ +\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,$, (12)
@@
4 Dirac Fermion in terms of Majorana fermions
@@
with each $\xi_{i}$ satisfying the Majorana condition [21]
@@
$\int_{\mathcal{M}}d^{3}x\,\bar{\psi}(i\cancel{\partial}-m)\psi=\int_{\mathcal{M}}d^{3}x\,\big[\bar{\xi}_{1}(i\cancel{\partial}-m)\xi_{1}+\bar{\xi}_{2}(i\cancel{\partial}-m)\xi_{2}\big]+\int_{\partial\mathcal{M}}d^{2}x\,\bar{\xi}_{1}\gamma^{\mu}\xi_{2}\,n_{\mu}\,$, (17)
@@
The boundary $\partial\mathcal{M}$ is $(1+1)$-dimensional, where chirality can be defined as in $(3+1)$ dimensions. Defining $\gamma^{5}=\gamma^{0}\gamma^{1}$}, which anticommutes with all $\gamma^{\mu}$, we can consider chiral boundary modes (eigenstates of $\gamma^{5}$) for the two Majorana flavours. These Majorana–Weyl fermions exist only in $(8k+2)$ spacetime dimensions [16]. On the boundary we define
@@
This result is consistent with the Abelian Chern–Simons description of a Dirac fermion at low energies, which carries $c_{-}=1$} [22, 15]. As we will emphasize later, the non-trivial boundary structure will play a key role in matching the edge modes across dualities.
Finally, we note that these boundary conditions emerge naturally from the variational principle used to derive the equations of motion. Although boundary conditions are often imposed externally, this approach allows the theory itself to determine consistent constraints [5, 4, 2].
5 Two Majorana flavours
@@
$\mathcal{L}_{\phi}[\vec{\phi}_{1},\vec{\phi}_{2},a]=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,$, (23)
@@
To analyze the phase structure, we follow the auxiliary-scalar method of [24]. Using the operator correspondence from [14, 1, 10],
@@
$\mathcal{L}_{\xi}[\xi_{1},\xi_{2}]-\varphi\sum_{i=1}^{2}\bar{\xi}_{i}\xi_{i}-v(\varphi)\;\longleftrightarrow\;\mathcal{L}_{\phi}[\vec{\phi}_{1},\vec{\phi}_{2},a]+\varphi\sum_{i=1}^{2}\vec{\phi}_{i}^{2}-v(\varphi)\,$, (26)
@@
The sign of $r$ determines the two possible regimes. For r>0{r>0}, $\varphi$ is heavy and can be integrated out, restoring the correspondence (24). For r<0{r<0}, $\varphi$ acquires a nonzero expectation value and induces mass terms for both fermions and scalars. On the bosonic side, the resulting quadratic term combines with $\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})$ to reproduce the full potential $V(\vec{\phi}_{1},\vec{\phi}_{2})$ in (13). Depending on the sign of $\langle\varphi\rangle\equiv\varphi_{0}$}, the two theories describe either a trivial insulator or a topological phase with chiral central charge $c_{-}=1$}. These phases coincide with those found in [24], as expected from the Dirac decomposition (17).
@@
$\sum_{i=1}^{2}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}$ $+N\text{CS}_{g}\;\longleftrightarrow\;|D_{b}z|^{2}-|z|^{4}+\text{CS}[b]_{1}\,$, (27)
@@
Again, the edge theories differ by one unit of chiral central charge ($c_{-}=1$} on the left, $c_{-}=0$} on the right), reconciled by the same boundary conditions preserving chirality. This transition connects two gapped phases with thermal Hall conductivities 0 and $\pm\tfrac{\pi}{6}T$ at low temperatures [11, 25].
It is illuminating to reinterpret these boundary conditions in the framework of (28). Following [14, 1], the Majorana fermion can be identified with the monopole operator of the gauge field, consistent with viewing this bosonization as a relativistic form of flux attachment. For two flavours, this follows from the equation of motion for the temporal component of the gauge field $a^{i,0}$ derived from (12):
@@
where $\mathscr{M}_{a}$ is the monopole operator of the gauge field [23]. The boundary condition (21) then translates to
@@
On the right-hand side of (28), the analogous identification involves the Dirac fermion [24]:
@@
6 Even number of Majorana fermions
@@
$\mathcal{L}_{\phi}[\vec{\phi}_{i},a]=\sum_{i=1}^{2N_{f}}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\ldots,\vec{\phi}_{2N_{f}})\;+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,$, (32)
@@
7 Ising–Gross–Neveu
Following Ref. [24], a theory of two Dirac fermions with staggered mass terms exhibits a transition between a trivial insulator and a quantum spin Hall insulator [6]. Once again, the use of Majorana variables provides a natural framework for constructing infrared dualities among purely bosonic theories.
@@ 7 Ising–Gross–Neveu
$\mathcal{L}_{\phi}[\vec{\phi}_{i},a,\tilde{a}]=\sum_{i=1}^{2}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}\\ +\sum_{i=3}^{4}|D_{\tilde{a}}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{3},\vec{\phi}_{4})+\text{CS}_{SO(\tilde{N})}[\tilde{a}]_{1}+(N+\tilde{N})\text{CS}_{g}\,.$ (35)
@@
and analogously for $SO(2)_{\tilde{\phi}}$ acting on $\tilde{\Phi}=(\vec{\phi}_{3},\vec{\phi}_{4})^{\top}$. These are genuine global symmetries, distinct from the $U(1)^{2}$ gauge symmetry acting on the $z_{i}$. On the $U(1)$ side, the corresponding topological symmetries are $U(1)_{T}^{2}$, under which monopole operators carry charge [23].
8 Pairing transition
Another transition discussed in Ref. [24] involves a superconducting deformation of the Dirac seed duality [12]: